# Engagement Maximization

### v5

@@ 4.3 Illustrative Examples
Figure 10: Credible beliefs with a decision-irrelevant $G$
5 Extensions
In this section, we consider several extensions to the model analyzed above. We first discuss the possibility that the agent does not fully attend to the principal’s signals, and show that this is never optimal for the agent. Second, we show that commitment to the signal structure plays a minor role in our model (in contrast to the importance of this assumption in the Bayesian persuasion literature). We next consider the case in which the cost of delay is increasing over time, representing either an increasing opportunity cost of time or diminishing utility from using the platform. We then discuss an alternative version of our model in which agents are not capacity constrained and the principal’s objective is to maximize the time until a decision is made.
5.1 Incentive Compatibility
In our main model, the agent solves a stopping problem given the principal’s chosen information structure. In practice, the agent may choose to process the principal’s chosen information structure in a selective way, which leads to an extra “incentive compatibility” requirement for the principal. We argue that the optimal policy is “incentive compatible” as is defined in Proposition 8.
Proposition 8.
Let $(P^{*},\tau^{*})$ be defined as in Theorem 1. For all stopping times $\hat{\tau}$ and all $\hat{\pi}\in\mathcal{P}(\mathcal{P}(X))$ that are a mean preserving contraction of $q_{\hat{\tau}}$, $\mathbb{E}^{\hat{\pi}}\left[\hat{\mu}(q)\right]-\mathbb{E}^{P^{*}}\left[\bar{\kappa}\hat{\tau}\big{|}\mathcal{F}_{0}\right]\leq\hat{\mu}(\bar{q}_{0})$.
Proof.
Observe that
$\mathbb{E}^{\hat{\pi}}\left[\hat{\mu}(q)\right]-\mathbb{E}^{P^{*}}\left[\bar{\kappa}\hat{\tau}\big{|}\mathcal{F}_{0}\right]\leq$ $\mathbb{E}^{P^{*}}\left[\hat{\mu}(q_{\hat{\tau}})-\bar{\kappa}\hat{\tau}\big{|}\mathcal{F}_{0}\right]$
$\leq$ $\mathbb{E}^{P^{*}}[\hat{\mu}(q_{\tau^{*}})-\bar{\kappa}\tau^{*}|\mathcal{F}_{0}]=\widehat{\mu}(\bar{q}_{0})$
The first inequality is from the convexity of $\widehat{\mu}$ and $\widehat{\pi}$ being a mean preserving contraction of $q_{\widehat{\tau}}$. The second inequality is the same as the main inequality of Proposition 2. The last equality is from the constraint in Equation R being binding for $\pi^{*}$. ∎
Any deviation of the agent leads to an alternative belief process $\widehat{q}_{t}$ and a corresponding stopping time $\widehat{\tau}$. Instead of modeling $\widehat{q}_{t}$ explicitly, observe that since $\widehat{q}_{t}$ is derived from a garbling of the signals that generate $q_{t}$, the law of $\widehat{q}_{\widehat{\tau}}$ (i.e. $\hat{\pi})$ is a mean preserving contraction of the law of $q_{\hat{\tau}}$. Proposition 8 therefore implies that the agent’s expected utility from such a deviation is weakly less than $\widehat{\mu}(\bar{q}_{0})$—the utility of fully attending the principal’s information.
The intuition of the result is simple: in the optimal policy, the principal is providing valuable information to the agent, to ensure that the agent does not stop using the platform. If the agent fails to attend to this information while continuing to use the platform, she incurs the cost of delay without fully benefiting from the information provided.
In the particular case in which the principal would like the agent to acquire decision-irrelevant information, the agent is indifferent between attending to all of the information provided by the principal and attending only to its decision-relevant components. Ignoring the decision-irrelevant information (i.e. ads) does not allow the agent to more rapidly acquire decision-relevant information; the arrival rate of decision-relevant information is determined by the principal, not the agent.
5.2 An Informed Principal
In this subsection, we modify our model by assuming that the principal knows $x\in X$ before choosing $P$. For the sake of brevity, we will discuss this case without providing a formal definition of equilibrium, and argue that our results continue to apply.
The key difference between this case and our main analysis is that the state $x\in X$ is now the principal’s “type,” and there is a set of conditional measures, $\{P_{x}\}_{x\in X}$, describing the belief processes the principal might induce in the agent conditional on each state $x\in X$. We will continue to think of these measures as being generated from signals the principal sends to the agent via Bayesian updating, and we assume that the agent observes the signal structures generating these signals. As a result, there is the potential for the principal to “signal via signal structures,” by selecting signal structures that signal the principal’s type, independently of the information conveyed by the signal structure itself.
To simplify our argument, we will assume that the agent observes, at date zero, the set of future signal structures the principal will offer. [^22] It should be clear from the argument that follows that this assumption is not essential. We claim that there is a pooling equilibrium in which all types of principal offer the signal structures that induce $P^{*}$, as in our main analysis, and the agent acts accordingly. The intuition is as follows: suppose a principal of type $x\in X$ deviated and offered a different set of signal structures. The agent would observe this deviation, and form off-path beliefs. Suppose that after such a deviation the agent places full support on a single type, $x^{\prime}\in X$. Given that the agent now believes she knows the true state, she will stop immediately, and therefore such a deviation is not beneficial for any principal.
Note that this argument holds even if the agent places the full support of her off-path beliefs on the type with the most reason to deviate (in the spirit of the D1 refinement). This point highlights a key difference between our model and standard Bayesian persuasion models. In standard Bayesian persuasion models, in which the principal cares about the action the agent ultimately chooses, an informed principal whose preferences are aligned with the agent has a strong incentive to reveal the true state, which gives refinements like D1 bite. In contrast, in our framework, no principal ever wants to reveal the true state, because each type of principal cares only about maximizing engagement and not about the action the agent ultimately chooses.
5.3 Increasing Costs of Delay
In this subsection, we consider the case in which the cost of delay is increasing over time. We denote the flow cost of delay at period $t$ by $\kappa(t).$ This case is of interest because it is natural to assume that the utility gained from using a platform declines, and the opportunity cost of time rises, the more time is spent on the platform. The same can be said for the utility of learning and the opportunity cost of time in the teacher-student context.
This case is not as tractable as the constant cost case. For this subsection only, we limit our analysis to the case of $H=G$. We also impose additional symmetry assumptions, not present in our main model, to allow us to prove our results. We should emphasize that we have no reason to believe, absent these symmetry assumptions, that our main results fail– but we have not been able to prove that they continue to hold, either. We will conjecture and verify that an equilibrium without commitment exists in which our characterization of optimal policy, suitably modified, continues to hold. Note that, that unlike the constant cost case, we will prove only the existence of an such an equilibrium in the absence of commitment, as opposed to proving the stronger claim that all equilibria share these properties.
Specifically, we will prove that there exists a $t_{0}\in\mathbb{R}_{+}$ such that if $q_{t_{0}}$ is uniform, then the optimal policy in the increasing cost case can be constructed from the optimal policies of the constant cost case.
Let $\pi^{*}(\kappa)$ be a solution to Proposition 1 under a constant cost of delay $\bar{\kappa}=\kappa$, given the uniform prior $\iota\in\mathcal{P}(X)$. The essence of our solution is that in each time $t$ the principal offers a belief process that either jumps to an element of the support of $\pi^{*}(\kappa(t))$ or remains constant. Define the stochastic process $\hat{q}_{s}$ as:
$\hat{q}_{s}=\iota+\mathbf{1}_{N_{\alpha}(s)\geq 1}\cdot(Q_{s}-\iota),$ (5)
where $Q_{s}\in\mathcal{P}(X)$ is a random variable distributed according to $\pi^{*}(\kappa(s))$ and $N_{\alpha}(s)$ is an independent Poisson counting process with an arbitrary parameter $\alpha>0$ and $N_{\alpha}(s)=0$ for $s\leq t_{0}$. Now define a time-changed version of this process,
$q_{t}=\hat{q}_{s(t)},$
with $s(t)=t$ for $t\in[0,t_{0}]$ and
$\frac{ds}{dt}=\frac{\chi}{\alpha E^{\pi^{*}(\kappa(t))}[H(q^{\prime})-H(\bar{q}_{0})]}$
for all $t>t_{0}$.
We will prove that if $q_{t_{0}}=\iota$, the law of this time-changed process from $t_{0}$ onward, $P^{*}$, along with $I_{t}^{H*}=\chi(t-t_{0})$ and the stopping rule $\tau^{*}=\inf\{t\in[t_{0},\infty):\hat{q}_{t}\neq\iota\}$, is part of a solution to
$\sup_{(P,I^{H},\tau)\in\mathcal{A}(\iota)}\mathbb{E}^{P}[H(q_{\tau})-H(\bar{q}_{0})|\mathcal{F}_{t_{0}}]$
subject to
[…]
(Recall here our assumption that $H=G$). That is, the principal-optimal policy $(P^{*},I^{H*},\tau^{*})$ can be implemented via this kind of signal structure.
Additional assumptions.
We introduce here a set of additional assumptions for the purpose of tractability. The first assumption states that the delay cost is non-decreasing, truncated at finite time, and high enough to guarantee interior solutions.
Assumption 2.
We assume that $\kappa(t):\mathbb{R}^{+}\to\mathbb{R}^{+}$ is non-decreasing, $\exists T$ s.t. $\kappa(t)$ is constant for $t\geq T$, and $\mathrm{Supp}(\pi^{*}(\kappa(0)))$ is in the interior of $\mathcal{P}(X)$.
The non-decreasing aspect of this assumption is natural, given that we would like to model an increasing opportunity cost of time or diminishing utility from using the platform. The existence of a $T$ after which the cost is constant is a technical device; our results will hold with $T$ arbitrarily large. The assumption that the optimal constant-cost policy is interior at time zero ensures that this remains true for all $t>0$, and simplifies our analysis. We view these assumptions as relatively innocuous.
Our substantive assumption is rotational symmetry. Let $R$ denote a matrix in $\mathbb{R}^{|X|^{2}}$. We say $R$ is a rotation if $Re_{x}\in\mathcal{P}(X)$, $R$ is full rank, stochastic, and satisfies $R^{J}=I$ for some $J\in\mathbb{N}^{+}$. A function $f:\mathcal{P}(X)\to\mathbb{R}$ is rotationally symmetric if $f(q)\equiv f(R(q))$ (for some rotation $R$). [^23] For example, any permutation matrix is a rotation. Therefore, any exchangeable function is rotationally symmetric.
Assumption 3.
$H$ and $\hat{\mu}$ are rotationally symmetric.
This kind of symmetry holds in our example, as the utility function and Shannon’s entropy are both symmetric about $q=\frac{1}{2}$.
Theorem 3.
Under Assumptions 2 and 3, there exists a $t_{0}<T$ s.t. if $q_{t_{0}}$ is uniform, then $(P^{*},I^{*},\tau^{*})$ is an optimal policy in the principal’s problem from $t_{0}$ onward.
Proof.
See the appendix, section C.1. ∎
The role of the initial time $t_{0}$ is analogous to the role of Assumption 1 in our main analysis. At $t=t_{0}$, the agent’s cost of delay is high enough to ensure that the incentive compatibility constraint is binding for the principal.
5.3.1 Example 1 (continued).
We illustrate Theorem 3 by continuing our example 1. We maintain the same model primitives except that $\kappa(t)$ no longer constant. The top panel of Figure 11 depicts $\kappa(t)$: $\kappa(t)$ strictly increasing from $[0,2]$ and stays constant afterwards. We choose $\kappa(t)$ take values close to $2$ for the purpose of facilitating comparisons with our previous examples (which had fixed $\bar{\kappa}=2$).
Figure 11: Increasing $\kappa(t)$ and the optimal policy.
The optimal policy is illustrated by the lower panel of Figure 11. The two black curves depict $\mathrm{Supp}(\pi^{*}(\kappa(t)))$. The optimal policy involves the same dilution structure as in Theorem 1, but with posterior beliefs varying over time. If one interprets how extreme the posterior beliefs are as the decision quality, then decision quality falls over time under increasing time costs.
5.4 Bounds on Belief Revision
In this subsection, we consider the case in which there is an exogenous bound on how much the agent can revise his posterior belief. Formally, for $d>0$, let $D^{d}$ denote the set of $\mathcal{P}(X)$-valued càdlàg functions whose jump size is bounded by $d$,
$D^{d}=\left\{f\in\Omega\big{|}\forall t\in\mathbb{R}_{+},\medspace|f(t)-f(t^{-})|\leq d\right\}.$
Let $\mathcal{A}^{d}(\bar{q}_{0})$ denote the subset of $\mathcal{A}(\bar{q}_{0})$ whose strategies are supported within $D^{d}$. Consider the optimization problem:
$J^{d}(\bar{q}_{0})$ $=\sup_{(P,I^{H},\tau)\in\mathcal{A}^{d}(\bar{q}_{0})}\mathbb{E}^{P}[G(q_{\tau})-G(\bar{q}_{0})|\mathcal{F}_{0}]$ (6)
s.t. $\tau\in\arg\max_{\tau^{\prime}\in\mathcal{T}}\mathbb{E}^{P}[\hat{u}(q_{\tau^{\prime}})-\bar{\kappa}\tau^{\prime}|\mathcal{F}_{0}].$
Equation 6 is the same as Definition 1 except that the set of admissible strategies is restricted to have no jumps of size greater than $d$.
We begin our analysis of this restricted problem by observing that the agent-optimal benchmark, $V^{B}$, is unchanged. This follows from results in Hébert and Woodford (2023), who show that the agent-optimal policy can be implemented by a pure diffusion process. Let $E^{A}=\left\{q\in\mathcal{P}(X)|V^{B}(q)>\widehat{u}(q)\right\}$ be the continuation region of the agent-optimal benchmark.
Now define the set of beliefs that can be reached by a jump of size no greater than $d$ from this continuation region:
$\bar{Q}^{d}=\left\{q\in\mathcal{P}(X)|\inf_{q^{\prime}\in E^{A}}|q-q^{\prime}|\leq d\right\}.$
Note that the definition of $\bar{Q}^{d}$ is independent of the prior $\bar{q}_{0}$.
As we argued previously, the beliefs $q_{t}$ must lie in the continuation region of the agent-optimal benchmark if $t<\tau$. Given that the maximum possible jump size is $d$, it follows that the stopping beliefs must lie in $\bar{Q}^{d}$.
Lemma 3.
If $(P,I^{H},\tau)$ is admissible in Equation 6, then $\mathrm{Supp}(q_{\tau})\subset\bar{Q}^{d}$.
Proof.
See the appendix, section C.2. ∎
Trivially, if the bound $d$ is sufficiently large, it does not restrict the principal, and the problem collapses to the problem considered in our main analysis.
Proposition 9.
$\lim_{d\to\infty}J^{d}(\bar{q}_{0})=J(\bar{q}_{0})$.
Proof.
For $d$ larger than the diameter of $\mathcal{P}(X)$, $\mathcal{A}^{d}(\bar{q}_{0})=\mathcal{A}(\bar{q}_{0})$. ∎
[…]
Proposition 10.
If $|X|=2$, then $J^{0}(\bar{q}_{0})$ is achieved by an agent-optimal policy.
Proof.
See the appendix, section C.3. ∎
When there are more than two states ($|X|>2$), the principal’s and agent’s optimal policies need not coincide. Consider as an example the case of two actions ($|A|=2$). The agent-optimal policy will in this case involve a diffusion on a line segment within the probability simplex (see Hébert and Woodford (2023)). The principal cannot induce the agent to follow this line segment beyond its endpoints, but can send the agent signals that cause the agent’s beliefs to move orthogonal to this line segment.
5.5 Optimal Policy without Capacity Constraints
In this subsection, we consider a modified version of our model with a constant cost of delay in which the agent has an unlimited capacity to acquire information, and the principal’s goal is to maximize the time spent by the agent on the platform. This modified model is the continuous-time analog of the discrete-time models studied by Knoepfle (2020) and Koh and Sanguanmoo (2022). The purpose of this subsection is illustrate the connection between these models and a special case of our more general framework. The agent-optimal policy in the modified model is for the agent to learn the optimal action with certainty immediately, as this avoids entirely the cost of delay.
@@ 5.5 Optimal Policy without Capacity Constraints
[…]
The first part of Lemma 1 remains applicable: if $\pi$ is the law of $q_{\tau}$, we must have $\mathbb{E}^{\pi}[\widehat{u}(q)-\hat{u}(\bar{q}_{0})]\geq\bar{\kappa}\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]$. Now observe that the expected utility under $\pi$ is bounded above by the utility of fully learning the state. Let $\pi^{\max}\in\mathcal{P}(X)$ be the unique probability measure that places full support on the extreme points of $\mathcal{P}(X)$ (i.e. the $e_{x}$ basis vectors), with $\pi^{max}(e_{x})=\bar{q}_{0,x}$.
By the convexity of $\hat{u}$, for all $\pi$ such that $E^{\pi}[q]=\bar{q}_{0}$,
$\mathbb{E}^{\pi^{max}}[\widehat{u}(q)-\hat{u}(\bar{q}_{0})]\geq\mathbb{E}^{\pi}[\widehat{u}(q)-\hat{u}(\bar{q}_{0})]\geq\bar{\kappa}\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]\geq\bar{\kappa}J(\bar{q}_{0}).$
It follows that $\bar{\kappa}^{-1}\mathbb{E}^{\pi^{max}}[\widehat{u}(q)-\hat{u}(\bar{q}_{0})]$ is an upper bound on the utility achievable in this problem.
But now observe that the $\alpha$-dilution of $\pi^{\max}$, as defined in (5), with intensity $\alpha=\bar{\kappa}(\mathbb{E}^{\pi^{max}}[\widehat{u}(q)-\hat{u}(\bar{q}_{0})])^{-1}$, achieves this bound. Moreover, the policy is incentive-compatible: at each instant, the agent compares the utility benefit of the signal’s arrival, $\alpha\mathbb{E}^{\pi^{max}}[\widehat{u}(q)-\hat{u}(\bar{q}_{0})]$, against the cost of delay, $\bar{\kappa}$, and is willing to continue. It follows that this policy is optimal.
This policy is the optimal policy in a special case of our main model. Consider the case of our main model in which $\bar{\kappa}=\chi$ and $G(q)=H(q)=\hat{u}(q)$. In this special case, $\pi^{\max}$ is an optimal policy in our relaxed problem (R), because the constraint in that problem satisfied for any policy, and the process described in Proposition 2 is exactly the process above. The intuition behind this equivalence is that in our main model, it is always optimal for the principal to exhaust the agent’s information processing capacity. If exhausting this capacity necessarily involves satisfying the incentive compatibility constraint, then the principal is free to choose the process that simultaneously exhausts the agent’s capacity and takes as long as possible to reach a decisive belief.
@@ A.1 Proof of Proposition 1
Assumption 1 implies that the constraint in (R) must be binding, as otherwise, the optimal $\pi^{*}$ will be fully revealing, leaving the agent with negative surplus. Then, (R) is a convex optimization problem that satisfies the conditions in Theorem 4 of Zhong [2018], which immediately the existence of $\lambda$ and a finite support solution $\pi^{*}$.
Next, we prove $\mathbb{E}^{\pi^{*}}[\bar{J}(q)]=0$. Suppose for the purpose of contradiction that $\mathbb{E}^{\pi^{*}}[J(q)]>0$. For each $q\in\text{Supp}(\pi^{*})$, let $\pi_{q}$ be the probability measure that implements $\bar{J}(q)$. Let $\pi^{\prime}=\mathbb{E}^{\pi^{*}(q)}[\pi^{q}]$. Then,
$\mathbb{E}^{\pi^{\prime}}[G(q)-G(\bar{q}_{0})]=$ $\mathbb{E}^{\pi^{*}(q_{1})}[\mathbb{E}^{\pi_{q_{1}}}[G(q)]-G(q_{1})+G(q_{1})-G(\bar{q}_{0})]$
$=$ $\mathbb{E}^{\pi^{*}(q_{1})}[\mathbb{E}^{\pi_{q_{1}}}[G(q)-G(q_{1})]]+\bar{J}(\bar{q}_{0})$
$=$ $\mathbb{E}^{\pi^{*}(q_{1})}[\bar{J}(q_{1})]+\bar{J}(\bar{q}_{0})$

### v6

@@ 4.3 Illustrative Examples
Figure 10: Credible interim beliefs with a decision-irrelevant $G$
5 Extension: Optimal Policy without Capacity Constraints
In this section, we consider a modified version of our model with a constant cost of delay in which the agent has an unlimited capacity to acquire information, and the principal’s goal is to maximize the time spent by the agent on the platform. This modified model is the continuous-time analog of the discrete-time models studied by jan2020dynamic and koh2022attention. The purpose of this section is to illustrate the connection between these models and a special case of our more general framework. The agent-optimal policy in the modified model is for the agent to learn the optimal action with certainty immediately, as this avoids entirely the cost of delay.
@@ 5 Extension: Optimal Policy without Capacity Constraints
$\tau\in\arg\max_{\tau^{\prime}\in\mathcal{T}}\mathbb{E}^{P}[\widehat{u}(q_{\tau^{\prime}})-\int_{0}^{\tau^{\prime}}\kappa dt|\mathcal{F}_{0}].$
The first part of Lemma 1 remains applicable: if $\pi$ is the law of $q_{\tau}$, we must have $\mathbb{E}^{\pi}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})]\geq\kappa\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]$. Now observe that the expected utility under $\pi$ is bounded above by the utility of fully learning the state. Let $\pi^{\max}\in\mathcal{P}(X)$ be the unique probability measure that places full support on the extreme points of $\mathcal{P}(X)$ (i.e. the $e_{x}$ basis vectors), with $\pi^{max}(e_{x})=\bar{q}_{0,x}$.
By the convexity of $\widehat{u}$, for all $\pi$ such that $E^{\pi}[q]=\bar{q}_{0}$,
$\mathbb{E}^{\pi^{max}}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})]\geq\mathbb{E}^{\pi}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})]\geq\kappa\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]\geq\kappa J(\bar{q}_{0}).$
It follows that $\kappa^{-1}\mathbb{E}^{\pi^{max}}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})]$ is an upper bound on the utility achievable in this problem.
But now observe that the $\alpha$-dilution of $\pi^{\max}$, as defined in (2), with intensity $\alpha=\kappa(\mathbb{E}^{\pi^{max}}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})])^{-1}$, achieves this bound. Moreover, the policy is incentive-compatible: at each instant, the agent compares the utility benefit of the signal’s arrival, $\alpha\mathbb{E}^{\pi^{max}}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})]$, against the cost of delay, $\kappa$, and is willing to continue. It follows that this policy is optimal.
This policy is the optimal policy in a special case of our main model. Consider the case of our main model in which $\kappa=\chi$ and $G(q)=H(q)=\widehat{u}(q)$. In this special case, $\pi^{\max}$ is an optimal policy in our relaxed problem (R), because the constraint in that problem satisfied for any policy, and the process described in Proposition 2 is exactly the process above. The intuition behind this equivalence is that in our main model, it is always optimal for the principal to exhaust the agent’s information processing capacity. If exhausting this capacity necessarily involves satisfying the incentive compatibility constraint, then the principal is free to choose the process that simultaneously exhausts the agent’s capacity and takes as long as possible to reach a decisive belief.
@@ A.1 Proof of Lemma 1
Condition (1): let $\tau^{\prime}\equiv 0$. Under this policy, the agent’s utility from stopping is $\widehat{u}(\bar{q}_{0})$. The optimality of $\tau$ and $\mathbb{E}^{\pi}[\hat{u}(q)]=\mathbb{E}^{P}[\hat{u}(q_{\tau})]$ implies condition (1).
Condition (2): This follows from $\mathbb{E}^{\pi}[H(q)]=\mathbb{E}^{P}[H(q_{\tau})]$, the optional stopping theorem, and the supermartingale property of $H(q_{t})-\chi t$. The last two imply that $\mathbb{E}^{P}[H(q_{\tau})-\chi\tau]\leq H(q_{0})$. ∎
A.2 Proof of Proposition 1
Proof.
Assumption 1 implies that the constraint in (R) must be binding, as otherwise, the optimal $\pi^{*}$ will achieve the principal’s first best, leaving the agent with negative surplus. Then, (R) is a convex optimization problem that satisfies the conditions in Theorem 4 of zhong2018information, which immediately implies the existence of $\lambda$ and of a finite support solution $\pi^{*}$.
Next, we prove $\mathbb{E}^{\pi^{*}}[\bar{J}(q)]=0$. Suppose for the purpose of contradiction that $\mathbb{E}^{\pi^{*}}[J(q)]>0$. For each $q\in\text{Supp}(\pi^{*})$, let $\pi^{q}$ be a probability measure that implements $\bar{J}(q)$. Define $\pi^{\prime}\in\Pi(\bar{q}_{0})$ by $d\pi^{\prime}(q^{\prime})=\mathbb{E}^{\pi^{*}(q)}[d\pi^{q}(q^{\prime})]$, observing that
$\mathbb{E}^{\pi^{\prime}(q^{\prime})}[q^{\prime}]=\mathbb{E}^{\pi^{*}(q)}[\mathbb{E}^{\pi^{q}(q^{\prime})}[q^{\prime}]]=\mathbb{E}^{\pi^{*}(q)}[q]=\bar{q}_{0}.$
Then,
$\mathbb{E}^{\pi^{\prime}(q^{\prime})}[G(q^{\prime})-G(\bar{q}_{0})]=$ $\mathbb{E}^{\pi^{*}(q)}[\mathbb{E}^{\pi^{q}(q^{\prime})}[G(q^{\prime})]-G(q)+G(q)-G(\bar{q}_{0})]$
$=$ $\mathbb{E}^{\pi^{*}(q)}[\mathbb{E}^{\pi^{q}(q^{\prime})}[G(q^{\prime})-G(q)]]+\bar{J}(\bar{q}_{0})$
$=$ $\mathbb{E}^{\pi^{*}(q)}[\bar{J}(q)]+\bar{J}(\bar{q}_{0})$